What is the energy in joules and eV of a photon in a radio wave from an AM station that has a 1500 kHz broadcast frequency?
------------ J
------------ eV
Answer #1

energy E =hf = 6.625x10^-34*1500x10^3 = 9.937 x10^-28 J

1eV = 1.6x10^-19 J

so

E = 9.937 x10^-28eV /1.6x10^-19 =6.2 x 10^-9 eV
Answer #2

E = h*frequency

= 6.6*10^-34*1500*1000

= 9.9* 10^-28 J

= 9.9*10^-9 eV